# Efficient Diffusion Models: A Survey

### v2

@@ 3.2.2 Adapter
Adapters are lightweight modules designed to enable efficient task adaptation by introducing small network layers into pre-trained models, allowing task-specific feature learning while keeping the original weights frozen. As illustrated in Figure 11, adapter layers are placed within the transformer block, positioned between normalization and feed-forward layers. Each adapter module consists of a down-projection, nonlinearity, and up-projection, which generates task-specific transformations without altering the core model’s structure.
This design significantly reduces memory and computational requirements, making adapters well-suited for tasks requiring lightweight parameter updates, such as text-to-image generation (T2I) and simulated domain adaptation (SimDA).
@@
Figure 11: Architecture of the Adapter module: demonstrating the integration of adapter layers within a transformer block to achieve efficient task adaptation by adding lightweight transformations, while keeping the core model weights frozen.
@@ 3.3 Efficient Sampling
The standard diffusion sampling process is computationally intensive because it requires sequentially executing a large number of denoising steps, with each step dependent on the output of the previous one as formulated in Eq(4). This inherent sequential dependency makes the sampling procedure time-consuming, resulting in significantly slower generation. To address these computational challenges, researchers have developed efficient sampling methods through four principal approaches. As illustrated in Figure 14, these encompass efficient SDE and ODE solvers that reduce the required number of function evaluations, advanced sampling scheduling strategies including parallel sampling techniques and timestep optimization methods, knowledge distillation techniques that transfer diffusion model capabilities to more efficient representations, and truncated sampling approaches that leverage early exit mechanisms and retrieval-based techniques to further accelerate the generation process while preserving output quality.
@@ 3.3.1 Efficient Solver
As we mentioned in Section. §2.1.3, although diffusion solvers avoid the need to compute the intractable partition function, they still require numerous function evaluations with fixed step sizes for SDEs in Eq(11) or generic numerical integration methods for ODEs in Eq.(12), resulting in slow sampling procedures and substantial computational costs. These limitations create substantial computational costs and lengthy generation times. In the following sections, we present advances in both SDE and ODE methods that address these efficiency constraints.
@@
ODE Solver. For efficiently implement the ODE-based solvers, Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a) builds upon the framework of Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), offering significant enhancements in sampling efficiency, which is one of the first models to leverage ODEs explicitly for the accelerating sampling process.
$q_{\sigma}(\mathbf{x}_{1:T}|\mathbf{x}_{0})=q_{\sigma}(\mathbf{x}_{T}|\mathbf{x}_{0})\prod_{t=2}^{T}q_{\sigma}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ (34)
Unlike DDPM’s Markovian forward process Eq.(1) where each state only depends on its immediate predecessor, DDIM utilizes the Non-Markovian Forward Process Eq.(34). These formulas allow each state not only to depend on its immediate predecessor but also on the initial state or a series of previous states. Specifically, Eq.(35) outlines how DDIM generates $\mathbf{x_{t-1}}$ from $\mathbf{x_{t}}$ by predicting the denoised observation, which essentially approximates reversing the diffusion process:
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})+\sigma_{t}\mathbf{\epsilon}_{t}$ (35)
@@
$\mathrm{d}\mathbf{x}(t)=\epsilon_{\theta}^{(t)}\left(\frac{\mathbf{x}(t)}{\sqrt{\sigma^{2}+1}}\right)\mathrm{d}\sigma(t)$ (36)
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\sigma_{d}\mathbf{F}_{\theta}(\frac{\mathbf{x}_{t}}{\sigma_{d}},c_{noise}(t))$ (37)
By introducing this simplified ODE parameterization, TrigFlow enables training large-scale CMs (up to 1.5B parameters) that achieve state-of-the-art performance with just two sampling steps, significantly reducing computational costs compared to DPM-solver (Lu et al., 2022) and other traditional diffusion models.
Moreover,  (Du et al., 2024) replaces SDE with ODE’s deterministic trajectory, combined with fixed noise and Consistency Distillation Sampling loss, which minimizes stochastic fluctuations and computational redundancy. These designs enable faster convergence while maintaining high fidelity.
@@ 3.3.2 Sampling Scheduling
Figure 15: Computation graph of Picard iterations, which introduces skip dependencies (Shih et al., 2024).
@@
ParaDiGMS (Shih et al., 2024) employs Picard iterations to parallelize the sampling process in diffusion models. This method models the denoising process using ordinary differential equations (ODEs) (Song et al., 2020c), where Picard iterations approximate the solution to these ODEs concurrently for multiple state updates. ParaDiGMS operates within a sliding window framework, enabling the simultaneous update of multiple state transitions. Each state is iteratively connected to different generations, allowing for information integration from several previous iterations Figure 15.
@@
Timestep Schedule. In the sampling process of diffusion models, the entire process is discrete, and the model progressively restores data from noise through a series of discrete timesteps. Each timestep represents a small denoising step that moves the model from its current state closer to the real data. The timestep schedule refers to the strategy for selecting and arranging these timesteps. It may involve distributing them evenly or performing denser sampling during key stages to ensure the efficiency of the sampling process and the quality of the generated results. Selecting an appropriate method to choose a series of timesteps can enable the sampling process to converge quickly, which the process is shown in Figure 16.
@@
However, optimizing an exact Evidence Lower Bound (ELBO) is typically not conducive to enhancing image quality. To address this, Align Your Steps (AYS) (Sabour et al., 2024) proposes a compute-efficient paradigm for diffusion models by co-optimizing sampling schedules and solvers. Through stochastic calculus-guided Kullback-Leibler Divergence Upper Bound (KLUB) minimization, as shown in  17, AYS derives adaptive schedules that reduce sampling steps by 40% while maintaining quality—achieving equivalent FID scores with 20 steps versus 30 steps in baseline methods. The optimized schedules are versatile across modalities and solver types, enabling plug-and-play acceleration without model retraining.
@@ 3.3.3 Truncated Sampling
The schematic in Figure 18 illustrates the time-dependent exit schedule used in ASE. As shown in the figure, the model progressively skips more building blocks as the diffusion process moves closer to the noise distribution, where the score estimation becomes easier. For example, in the DiT model (Peebles & Xie, 2023), the blocks are dropped progressively as the timestep approaches 1, closer to the noise regime. Conversely, when the timestep is closer to the data regime, more blocks are retained to ensure accurate score estimation. Moreover, the U-ViT model (Bao et al., 2023) follows a similar dropping schedule, but due to the long skip connections between the encoder and decoder, only decoder blocks are skipped. This preserves critical information from the encoder while still speeding up the process.
@@
Through carefully designed time-varying exit schedules, the method significantly accelerates sampling speed while preserving generation quality. It can also be seamlessly integrated with various diffusion model solvers to enhance overall efficiency further.
@@ 3.4.1 Quantization
Post-Training Quantization. PTQ involves selecting operations for quantization, collecting calibration samples, and determining quantization parameters for weights and activations. While collecting calibration samples is straightforward for CNNs and ViTs using real training data, it poses a challenge for Diffusion Models (DMs). In DMs, the inputs are generated samples $\mathbf{x}_{t}$ at various time steps (t = 0, 1, …, T), where T is large to ensure convergence to an isotropic Normal distribution. To address this issue, Shang et al. (2023) proposes PTQ4DM, the first DM-specific calibration set collection method, generating calibration data across all time steps with a specific distribution. However, their explorations remain confined to lower resolutions and 8-bit precision. Q-Diffsuion (Li et al., 2023) propose a time step-aware calibration data sampling to improve calibration quality and apply BRECQ (Li et al., 2021), which is a commonly utilized PTQ framework, to improve performance. Furthermore, compared to conventional PTQ calibration methods, they identify the accumulation of quantization error across time steps as another challenge in quantizing DMs Figure 25 (a). Therefore, they also propose a specialized quantizer for the noise estimation network shown in Figure 25 (b). Based on Q-Diffusion, Kim et al. (2024b) find that inaccurate computation during the early stage of the reverse diffusion process has minimal impact on the quality of generated images. Therefore, they introduce a method that focuses on further reducing the number of activation bits for the early reverse diffusion process while maintaining high-bit activations for the later stages. Lastly, He et al. (2024) presents PTQD, a unified formulation for quantization noise and diffusion perturbed noise. Additionally, they introduce a step-aware mixed precision scheme, which dynamically selects the appropriate bitwidths for synonymous steps.
@@
So et al. (2024) proposes a novel quantization method that enhances output quality by dynamically adjusting the quantization interval based on time step information.
Figure 27: Illustrations of the pruning.
The proposed approach integrates with the Learned Step Size Quantization (Esser et al., 2019) framework, replacing the static quantization interval with a dynamically generated output from the Time-Dynamic Quantization module. This dynamic adjustment leads to significant improvements in the quality of the quantized outputs. He et al. (2023) introduces a quantization-aware low-rank adapter that integrates with model weights and is jointly quantized to a low bit-width. This approach distills the denoising capabilities of full-precision models into their quantized versions, utilizing only a few trainable quantization scales per layer and eliminating the need for training data.
@@ 4.2 Parallel Computing
Parallel computing Figure 30 plays a critical role in the efficient execution of diffusion models, especially given the computation-intensive nature of these algorithms. Recent advances in parallel computing strategies have enabled significant improvements in inference speed and scalability, often without compromising the quality of the generated output (Li et al., 2024a; Wang et al., 2024a; Li et al., 2024b; Tian et al., 2024). This section highlights several notable contributions that tackle the challenge of parallelizing diffusion models across multiple GPUs and other distributed architectures.
@@ 6 Future Work
Despite the significant progress made in efficient diffusion models, several promising research directions remain open. In particular, we identify the following key areas for future work:

### v3

@@ 3.2.2 Adapter
Adapters are lightweight modules designed to enable efficient task adaptation by introducing small network layers into pre-trained models, allowing task-specific feature learning while keeping the original weights frozen. As illustrated in Figure 11, adapter layers are placed within the transformer block, positioned between normalization and feed-forward layers. Each adapter module consists of a down-projection, nonlinearity, and up-projection, which generates task-specific transformations without altering the core model’s structure. This design significantly reduces memory and computational requirements, making adapters well-suited for tasks requiring lightweight parameter updates, such as text-to-image generation (T2I) and simulated domain adaptation (SimDA).
@@ 3.3 Efficient Sampling
The standard diffusion sampling process is computationally intensive because it requires sequentially executing a large number of denoising steps, with each step dependent on the output of the previous one as formulated in Eq.(4). This inherent sequential dependency makes the sampling procedure time-consuming, resulting in significantly slower generation. To address these computational challenges, researchers have developed efficient sampling methods through four principal approaches. As illustrated in Figure 14, these encompass efficient SDE and ODE solvers that reduce the required number of function evaluations, advanced sampling scheduling strategies including parallel sampling techniques and timestep optimization methods, knowledge distillation techniques that transfer diffusion model capabilities to more efficient representations, and truncated sampling approaches that leverage early exit mechanisms and retrieval-based techniques to further accelerate the generation process while preserving output quality.
@@ 3.3.1 Efficient Solver
As mentioned in §2.1.3, although diffusion solvers avoid the need to compute the intractable partition function, they still require numerous function evaluations with fixed step sizes for SDEs in Eq.(11) or generic numerical integration methods for ODEs in Eq.(12), resulting in slow sampling procedures and substantial computational costs. These limitations create substantial computational costs and lengthy generation times. In the following sections, we present advances in both SDE and ODE methods that address these efficiency constraints.
@@
ODE Solver. For efficiently implement the ODE-based solvers, Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a) builds upon the framework of Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), offering significant enhancements in sampling efficiency, which is one of the first models to leverage ODEs explicitly for the accelerating sampling process. Unlike DDPM’s Markovian forward process in Eq.(1) where each state only depends on its immediate predecessor, DDIM utilizes the Non-Markovian Forward Process in Eq.(34):
$q_{\sigma}(\mathbf{x}_{1:T}|\mathbf{x}_{0})=q_{\sigma}(\mathbf{x}_{T}|\mathbf{x}_{0})\prod_{t=2}^{T}q_{\sigma}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}).$ (34)
These formulas allow each state not only to depend on its immediate predecessor but also on the initial state or a series of previous states. Specifically, Eq.(35) outlines how DDIM generates $\mathbf{x_{t-1}}$ from $\mathbf{x_{t}}$ by predicting the denoised observation, which essentially approximates reversing the diffusion process:
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})+\sigma_{t}\mathbf{\epsilon}_{t}.$ (35)
@@
$\mathrm{d}\mathbf{x}(t)=\epsilon_{\theta}^{(t)}\left(\frac{\mathbf{x}(t)}{\sqrt{\sigma^{2}+1}}\right)\mathrm{d}\sigma(t).$ (36)
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\sigma_{d}\mathbf{F}_{\theta}(\frac{\mathbf{x}_{t}}{\sigma_{d}},c_{noise}(t)).$ (37)
By introducing this simplified ODE parameterization, TrigFlow enables training large-scale CMs (up to 1.5B parameters) that achieve state-of-the-art performance with just two sampling steps, significantly reducing computational costs compared to DPM-solver (Lu et al., 2022) and other traditional diffusion models. Moreover, (Du et al., 2024) replaces SDE with ODE’s deterministic trajectory, combined with fixed noise and Consistency Distillation Sampling loss, which minimizes stochastic fluctuations and computational redundancy. These designs enable faster convergence while maintaining high fidelity.
@@ 3.3.2 Sampling Scheduling
ParaDiGMS (Shih et al., 2024) employs Picard iterations to parallelize the sampling process in diffusion models. This method models the denoising process using ordinary differential equations (ODEs) (Song et al., 2020c), where Picard iterations approximate the solution to these ODEs concurrently for multiple state updates. ParaDiGMS operates within a sliding window framework, enabling the simultaneous update of multiple state transitions. Each state is iteratively connected to different generations, allowing for information integration from several previous iterations.
@@
Timestep Schedule. In the sampling process of diffusion models, the entire process is discrete, and the model progressively restores data from noise through a series of discrete timesteps. Each timestep represents a small denoising step that moves the model from its current state closer to the real data. The timestep schedule refers to the strategy for selecting and arranging these timesteps. It may involve distributing them evenly or performing denser sampling during key stages to ensure the efficiency of the sampling process and the quality of the generated results. Selecting an appropriate method to choose a series of timesteps can enable the sampling process to converge quickly, as shown in Figure 15.
@@
However, optimizing an exact Evidence Lower Bound (ELBO) is typically not conducive to enhancing image quality. To address this, Align Your Steps (AYS) (Sabour et al., 2024) proposes a compute-efficient paradigm for diffusion models by co-optimizing sampling schedules and solvers. Through stochastic calculus-guided Kullback-Leibler Divergence Upper Bound (KLUB) minimization, as shown in Figure 16, AYS derives adaptive schedules that reduce sampling steps by 40% while maintaining quality—achieving equivalent FID scores with 20 steps versus 30 steps in baseline methods. The optimized schedules are versatile across modalities and solver types, enabling plug-and-play acceleration without model retraining.
@@ 3.3.3 Truncated Sampling
The schematic in Figure 17 illustrates the time-dependent exit schedule used in ASE. As shown in the figure, the model progressively skips more building blocks as the diffusion process moves closer to the noise distribution, where the score estimation becomes easier. For example, in the DiT model (Peebles & Xie, 2023), the blocks are dropped progressively as the timestep approaches 1, closer to the noise regime.
@@
Conversely, when the timestep is closer to the data regime, more blocks are retained to ensure accurate score estimation. Moreover, the U-ViT model (Bao et al., 2023) follows a similar dropping schedule, but due to the long skip connections between the encoder and decoder, only decoder blocks are skipped. This preserves critical information from the encoder while still speeding up the process. Through carefully designed time-varying exit schedules, the method significantly accelerates sampling speed while preserving generation quality. It can also be seamlessly integrated with various diffusion model solvers to enhance overall efficiency further.
@@ 3.4.1 Quantization
Post-Training Quantization. PTQ involves selecting operations for quantization, collecting calibration samples, and determining quantization parameters for weights and activations. While collecting calibration samples is straightforward for CNNs and ViTs using real training data, it poses a challenge for Diffusion Models (DMs). In DMs, the inputs are generated samples $\mathbf{x}_{t}$ at various time steps (t = 0, 1, …, T), where T is large to ensure convergence to an isotropic Normal distribution. To address this issue, Shang et al. (2023) proposes PTQ4DM, the first DM-specific calibration set collection method, generating calibration data across all time steps with a specific distribution. However, their explorations remain confined to lower resolutions and 8-bit precision. Q-Diffsuion (Li et al., 2023) propose a time step-aware calibration data sampling to improve calibration quality and apply BRECQ (Li et al., 2021), which is a commonly utilized PTQ framework, to improve performance. Furthermore, compared to conventional PTQ calibration methods, they identify the accumulation of quantization error across time steps as another challenge in quantizing DMs (Figure 23 (a)). Therefore, they also propose a specialized quantizer for the noise estimation network shown in Figure 23 (b). Based on Q-Diffusion, Kim et al. (2024b) find that inaccurate computation during the early stage of the reverse diffusion process has minimal impact on the quality of generated images. Therefore, they introduce a method that focuses on further reducing the number of activation bits for the early reverse diffusion process while maintaining high-bit activations for the later stages. Lastly, He et al. (2024) presents PTQD, a unified formulation for quantization noise and diffusion perturbed noise. Additionally, they introduce a step-aware mixed precision scheme, which dynamically selects the appropriate bitwidths for synonymous steps.
@@
So et al. (2024) proposes a novel quantization method that enhances output quality by dynamically adjusting the quantization interval based on time step information. The proposed approach integrates with the Learned Step Size Quantization (Esser et al., 2019) framework, replacing the static quantization interval with a dynamically generated output from the Time-Dynamic Quantization module. This dynamic adjustment leads to significant improvements in the quality of the quantized outputs. He et al. (2023) introduces a quantization-aware low-rank adapter that integrates with model weights and is jointly quantized to a low bit-width. This approach distills the denoising capabilities of full-precision models into their quantized versions, utilizing only a few trainable quantization scales per layer and eliminating the need for training data.
@@ 4.2 Parallel Computing
Parallel computing (Figure 27) plays a critical role in the efficient execution of diffusion models, especially given the computation-intensive nature of these algorithms. Recent advances in parallel computing strategies have enabled significant improvements in inference speed and scalability, often without compromising the quality of the generated output (Li et al., 2024a; Wang et al., 2024a; Li et al., 2024b; Tian et al., 2024). This section highlights several notable contributions that tackle the challenge of parallelizing diffusion models across multiple GPUs and other distributed architectures.
@@ 6 Future Work
Despite significant progress made in efficient diffusion models, several promising research directions remain open. In particular, we identify the following key areas for future work.
@@ 7 Conclusion
8 Acknowledgement
We would like to thank the action editor Ming-Hsuan Yang and anonymous reviewers of Transactions on Machine Learning Research for their helpful and constructive comments.